D. From a location 4.0 m away, you listen to a jackhammer breaking concrete. 1. How far away would you have to stand in orderthis is one question with 3 parts
Answer #1

intensity=\frac{energy}{time \times area}

intensity\propto \frac{1}{r^2}

\frac{I_1}{I_2}=\frac{r_2^2}{r_1^2}

\frac{I_1}{\frac{I_1}{8}}=\frac{r_2^2}{(4)^2}

\frac{8\times I_1}{I_1}=\frac{r_2^2}{(4)^2}

{r_2^2}=8\times (4)^2

r_2=11.31 m

(2) I\propto A^2

\frac{I_1}{I_2}=\frac{A_1^2}{A_2^2}

8=\frac{A_1^2}{A_2^2}

A_2=\frac{A_1}{\sqrt{8}}

(3)\beta(dB)=10log(\frac{I_2}{I_1})

\beta(dB)=10 log(\frac{1}{8})

\beta(dB)=-9.03